All open issues [^] (1 - 10 / 10) |
0000199
|
System temperature at L-band is way above what is expected
front-end - 06-03-25 17:04
|
0000109
|
The TP of DX(MFFE-13) is unstable.
0.unknown/unspecified - 13-11-24 03:25
|
0000048

|
Airco failure
0.unknown/unspecified - 14-09-24 11:38
|
0000273
|
The TP of 3X is unstable in all bands.
back-end - 24-01-12 13:59
|
0000201
|
Spikes on several baselines (RFI)
0.unknown/unspecified - 04-01-11 14:33
|
0000126
|
RFI in 7X7X in line obs with 1 band, 1024ch, 2polars,BW=10MHz
0.unknown/unspecified - 04-01-11 14:32
|
0000200
|
Peaks in the crosscorrelations in L-band observations
0.unknown/unspecified - 10-02-10 11:13
|
0000113
|
The TP of AX(MFFE-02) is unstable at Lband.
0.unknown/unspecified - 06-04-09 07:14
|
0000108
|
The TPs of 4X,4Y(MFFE-16) are unstable.
0.unknown/unspecified - 18-03-09 08:41
|
0000045
|
Failed observation due to resetting MFFEs
0.unknown/unspecified - 08-09-08 08:06
|